OpenStudy (anonymous):

Solve ln(x) + ln(x - 1) = 1 Please explain steps

OpenStudy (anonymous):

Using properties of logarithims: ln(x) + ln(x - 1) = 1 ln(x(x-1))=1 e^ln(x(x-1))=e^1 x(x-1)=e x^2-x-e=0 Using quadratic equation x=(1+sqrt(1+4 e))/2. Note the negative solution is not in the domain of lnx of ln(x-1) so it's not an answer